OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now? Numerical Answers Expected!

OpenStudy (lollylau):

13

OpenStudy (anonymous):

Let Coach Johnson be C Time be T

OpenStudy (anonymous):

C = coach age T = Tim age C = T + 10 C-3 = 2(T-3), so C = 2T - 6 + 3 = 2T - 3 2T - 3 = T+ 10 T = 13, Tim's age now so coach is 23 three years ago they were 10 and 20